faucet a, b, and c can fill out the pool in order in 6, 8, 12 hours. If we keep them on 3 hours there will be 3000 liters amount of water. What is the capacity of the pool?
Answer by stanbon(75887)

faucet a, b, and c can fill out the pool in order in 6, 8, 12 hours. If we keep them on 3 hours there will be 3000 liters amount of water. What is the capacity of the pool?
----------
rate of a: 1/6 job/hr
rate of b: 1/8 job/hr
rate of c: 1/12 job/hr
-----
Note: time*rate = work done
==================================
Total Rate:
(1/6 + 1/8 + 1/12) = (4+3+2)/24 = 9/24 = 3/8 job/hr
---
Let the capacity of the pool be "x" liters
Equation:
(3/8)x = 3000
(1/8)x = 1000
x = 8000 liters (pool capacity)
==================================
Cheers,
Stan H.